OpenStudy (casadie1004):

Please help, will medal and fan!!! For the following system of equations what is the x value of the solution? -x+2y=6 6y=x+18

OpenStudy (anonymous):

ok that was wrong sorry

OpenStudy (anonymous):

you can write the first equation as \[x=2y-6\] substitute that in to the second and solve \[6y=2y-6+18\] for \(y\)

OpenStudy (anonymous):

did that help or no?

OpenStudy (casadie1004):

3?

OpenStudy (anonymous):

yes, \(y=3\) now you need to find \(x\)

OpenStudy (casadie1004):

the question only asked for y, but thank u.

OpenStudy (anonymous):

actually it asked for x

OpenStudy (anonymous):

i mean that is what you wrote anyways

OpenStudy (casadie1004):

lol nvm ur right i was reading off of the wrong question on the test. how do i find x?

OpenStudy (anonymous):

you know that \[-x+2y=6\] and that \(y=3\) so you can solve \[-x+6=6\]for \(x\)

OpenStudy (anonymous):

or, since \[6y=x+18\] you can solve \[18+x=18\] either way

OpenStudy (casadie1004):

0?

OpenStudy (anonymous):

yes, both times, which means it is certainly correct

OpenStudy (casadie1004):

Thank you! :)

OpenStudy (casadie1004):

Can you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (casadie1004):

solve the following system of equations using substitution. what is the value of y. 2x+3y=105 x+2y=65

OpenStudy (casadie1004):

a.15 b.25 c.40 d.65

OpenStudy (anonymous):

here you can write the second one as \[x=65-2y\] and substitute in to the first to solve \[2(65-2y)+3y=105\]

OpenStudy (anonymous):

sorry got logged out can you solve that one?

OpenStudy (casadie1004):

yeah sorry i walked away from the computer, 1 second. almost done

OpenStudy (casadie1004):

y=25?

OpenStudy (anonymous):

oui

OpenStudy (casadie1004):

what?

OpenStudy (anonymous):

yes

OpenStudy (casadie1004):

thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

that it?

OpenStudy (casadie1004):

yes :)

OpenStudy (anonymous):

whew